OpenStudy (anonymous):

This one i am really stuck with (3x-2)^3-27

OpenStudy (anonymous):

what are you supposed to do?

OpenStudy (anonymous):

sorry factor

OpenStudy (anonymous):

This is a difference of two cubes. We can see this because: \[(3x-2)^3-27 \Leftrightarrow (3x-2)^3-3^3\]

OpenStudy (anonymous):

so you want to use the formula: \[a^3-b^3 = (a-b)(a^2+ab+b^2), a = 3x-2, b = 3\]

OpenStudy (id21):

\[(3x-2)^3-27=(3x-2)^3-3^3=(3x-2-3)((3x-2)^2+3(3x-2)+3^2)\] etc.

OpenStudy (anonymous):

ok, i used the difference of cubes and now i have (3x-2)(9x^2+6x+4)-27 how do i get the 27 out of there

OpenStudy (anonymous):

when you use the difference of two cubes, you should have any terms sticking out like that. that -27 probably doesnt belong there.

OpenStudy (anonymous):

shouldnt*

OpenStudy (anonymous):

ok i see what i'm doing wrong